# Variational Bayesian Bow tie Neural Networks with Shrinkage

### v2

@@ 3.3 Inferring the Network Structure
The choice of the network architecture has significant practical implications on the generalization of the model, and so sparsity-promoting priors for the network weights have emerged as promising approach to allow for a data-driven choice of architecture. Instead of the classical two-group discrete mixture priors, in this article, we focus on a class of continuous shrinkage priors. While this results in more tractable computations, it also implies non-zero posterior means and draws and doesn’t lead to automatic network architecture selection. Several post-processing methods have been proposed to yield a sparse solution (see e.g.Piironen et al. (2020); Li and Pati (2017); Griffin (2024) ). The method known as decoupling shrinkage and selection (DSS) Hahn and Carvalho (2015) obtains sparse estimates of the weights by minimizing the sum of the predictive loss function with a parsimony-inducing penalty. An alternative approach is the the penalized credible regions (PenCR) method Zhang et al. (2021), which identifies the "sparsest" solution in posterior credible regions corresponding to different levels; it is shown to perform well in the case of global-local shrinkage priors and under certain assumptions, PenCR produces the same results as DSS. Similarly, we propose to make use of credible intervals to select nodes. Following Li and Lin (2010), we implement an automatic credible interval criterion which selects a node as long as its credible interval doesn’t cover zero. Specifically, recall that the variational posterior of the weights is $W_{l,d,d^{\prime}}\sim\text{N}(m^{W}_{l,d,d^{\prime}},B^{W}_{l,d,d^{\prime}})$ for $l=1,\ldots,L,d=1,\dots,D_{l},d^{\prime}=1,\ldots,D_{l-1}$, where $\mathbf{m}^{W}_{l,d}$ and $\mathbf{B}^{W}_{l,d}$ denote the subsets of the mean $\mathbf{m}_{l,d}$ and covariance matrix $\mathbf{B}_{l,d}$ corresponding to the weights. Then, we obtain sparse weights $\widehat{W}_{l,d,d^{\prime}}$ with sparse variational distribution $\widehat{q}(b_{l,d},\widehat{\mathbf{W}}_{l,d})$ for some $l\in\mathcal{L}\subseteq\{1,\ldots,L+1\},\;d\in\mathcal{D}_{l}\subseteq\{1,\ldots,D_{l}\},d^{\prime}\in\mathcal{D}_{l-1}\subseteq\{1,\ldots,D_{l-1}\}$, defined by setting:
@@
where $Q(W_{l,d,d^{\prime}}<0)=1-Q(W_{l,d,d^{\prime}}>0)=\Phi(-m^{W}_{l,d,d^{\prime}}/\sqrt{(B^{W}_{l,d})_{d^{\prime}d^{\prime}}})$.
The threshold $\kappa$ is chosen to control the Bayesian false discovery rate, which is calculated as
@@
Once we sweep through $\mathcal{Q}_{l,d.d^{\prime}}$, we do a backwards pass to remove the nodes with no connections. If the node has no outgoing connections then all the incoming connections need to be removed, and conversely, if the node has no incoming connections, then all the outgoing connections can be removed. An example of the network resulting after applying the Algorithm 2 is illustrated by the Figure 7.
@@ 3.4 Predictions
Thus, before computing predictions, we first iterate to update the variational predictive distributions of $\mathbf{a}_{*}$, $\bm{\gamma}_{*}$, and $\bm{\omega}_{*}$. The corresponding ELBO (dervived in the Section B.2) is monitored for convergence. While a closed-form expression for the integral in Equation 15 is unavailable, generating samples from the variational predictive is straightforward;
@@
We can also compute the expectation and variance of $\mathbf{y}_{*}$ in closed-form. Specifically, the expectation of $\mathbf{y}_{*}$ under the variational predictive distribution is:
@@
Sparse prediction.
To save on both computation and storage, the variational predictive distribution can be computed based on the sparse variational posterior (Section 3.3). For a new data point $\mathbf{x}_{*}$, we obtain expectation and variance of $\mathbf{y}_{*}$ by first computing the sparse versions of variational predictive distributions $\widehat{q}(\mathbf{a}_{*}),\widehat{q}(\bm{\gamma}_{*})$ and $\widehat{q}(\bm{\omega}_{*})$ as in Equations 17, 14 and 16 by plugging $\widehat{q}(b_{l,d},\widehat{\mathbf{W}}_{l,d})$ instead of the $q(b_{l,d},\mathbf{W}_{l,d})$, which only requires updates for the subset of nodes with nonzero weights.
@@ 3.5 Ensembles of Variational Approximations
While the variational algorithm described in Section 3.1 increases the ELBO at each epoch, the ELBO is a non-convex function of the variational parameters and only convergence to a local optimum is guaranteed. Due to identifiability issues, the posterior distribution of a Bayesian neural network is highly multimodal, and exploring this posterior is notoriously challenging Papamarkou et al. (2022). A single variational approximation tends to concentrate around one mode and can understate posterior uncertainty. Several approaches have been proposed to overcome such issues. Recently, Ohn and Lin (2024) introduce adaptive variational inference which achieves optimal posterior contraction rate and model selection consistency by considering several variational approximations obtained in different models. Yao et al. (2022) introduce an approach which uses parallel runs of inference algorithms to cover as many modes of the posterior distribution as possible and then combines these using Bayesian stacking.
In a similar but simpler fashion to the above proposal, we consider an ensemble of variational approximations, obtained by running in parallel the variational algorithm multiple times with different random starting points. In this case, letting $k=1,\ldots,K$ index the different variational approximations, we compute the weight $w_{k}$ associated with each approximation in accordance with the optimization objective, the ELBO:
$w_{k}\propto\exp\left(\text{ELBO}_{k}\right).$
We can interpret this as a Bayesian model averaging across the $K$ different models/approximations. While in an ideal setting, the weights would be proportional to the marginal likelihood for each model, the use of the ELBO is motivated as it provides a lower bound to the marginal likelihood and can be computed in closed form. Next, we compute predictions by taking a weighted average of the predictive distributions of each model (given in Equation 19), that is
@@ 4 Experiments
We evaluate the variational bow tie neural network (VBNN) on several datasets. First, we consider a simple nonlinear synthetic example to compare with a ground truth. We then validate VBNN on the diabetes dataset, first considered in Efron et al. (2004) to demonstrate the least angle regression (LARS) algorithm for variable selection, and subsequently, used in different proposals for sparsity-promoting priors and algorithms (e.g. Park and Casella (2008); Li and Lin (2010)). Lastly, we consider a range of popular regression datasets from the UCI Machine Learning Repository M. et al. (2007).
The importance of suitable initialization choice in NNs is well known Wenzel et al. (2020); Daniely et al. (2016); He et al. (2015), and we design two possible random initialization schemes of the VBNN, which are described in Section C.1 and used in all experiments. Convergence of the ELBO is monitored during the training and prediction stages, where if three consecutive measurements of ELBO for training differ by less than the specified threshold, the phase is stopped and the model moves to the prediction stage, where we proceed similarly. In most experiments, the thresholds during the training and prediction stages are set, respectively, to $51e-$ 5 and $41e-$ 4. We compare the performance of VBNN to two popular variational frameworks under the mean-field assumption: the stochastic variational inference approximation (SVI) implemented in Numpyro Phan et al. (2019), and Bayes by Backprop (BBB) Blundell et al. (2015) implemented with Pytorch. For all the datasets, we evaluate the performance over 10 random splits, where we use $90\%$ of the data for the training and $10\%$ for testing the model. We normalize the input but do not re-scale the output, we record the root mean squared error (RMSE), the predictive negative log-likelihood of the test data (NLL) and the empirical coverage, see Appendix C for additional implementation details and the definition of each metric.
@@ 4.1 Simulated example
We construct a synthetic dataset generated by first uniformly sampling a two-dimensional input vector $\mathbf{x}_{n}=(x_{n,1},x_{n,2})$, with $x_{n,d}\sim\text{Unif}([-2,2])$, and assume only the first feature influences the output: $y_{n}=f(x_{n,1})+\epsilon_{n}=0.1x_{n,1}^{2}+10\sin(x_{n,1})+\epsilon_{n}$, where $\epsilon_{n}\sim\text{N}(0,0.5)$. Then, the dataset consisting of $N=300$ observations is used to investigate the performance of VBNN compared to the SVI and BBB baselines as we increase the number of hidden layers, setting $L=1,2$ or $5$, whilst keeping the number of hidden units in each layer fixed to $D_{H}=20$.
In general, for this simple non-linear example, the performance tends to deteriorate with increasing architecture complexity (larger depth), and the VBNN is the most robust to this choice (see Figure 5, which compares the performance of three of the models as a function of depth). In addition, as the number of hidden layers increases, SVI and BBB provide overly wide credible intervals while VBNN more closely aligns with the desired coverage (see Figure 6, which illustrates the empirical coverages for the observations).
Figure 5: Performance on the simulated dataset as the depth increases. VBNN is more robust to the choice of depth and overparameterization.
Figure 6: Empirical coverage for the observations for the simulated dataset for three different settings of network’s depth.
@@
For each depth, Figure 7 illustrates the predictive means and uncertainties computed for the observations as well as DAGs of networks’ structures obtained after the post-process node selection algorithm described in Section 3.3, where the Bayesian false discovery rate is constrained by setting the error rate to $\alpha=0.01$. The sparsity-promoting prior combined with the node selection algorithm can effectively prune the over-parametrized neural networks; for example, the sparse one-layer neural network contains only $11$ hidden nodes with $30$ total edges/weights from the initial $D_{H}=20$ with $60$ total edges/weights. Moreover, the estimated regression function and credible intervals both from the variational predictive and the sparse variation predictive distribution, recover the true function well. In this way, VBNN provides an effective tool to reduce predictive computational complexity and storage as well as ease interpretation. Note that the predictions show no relation with the coordinate $x_{2}$ (Figure 7(d)), recovering the true function, but some of the connections from $x_{2}$ are still present in the sparse network, due to identifiability issues, although with overall low weight (Figure 7(e)).
(a) $L=1$
(b) $L=2$
(c) $L=5$
(d) The predictive mean as a function of the second coordinate for $L=1$.
(e) Predictive means of the weights (on the left) and of the sparse weights (on the right) for $L=1$.
@@
Figure 7: Predictive means and uncertainty estimates computed for the observations and the architecture of the network for the bound on the FDR $\alpha=0.01$ for different settings of network’s depth.
@@ 4.2 Diabetes example
The diabetes data consists of $n=442$ entries obtained for $p=10$ input variables, the response is a quantitative measurement of disease progression. The predictors are age, sex, body mass index, average blood pressure and six blood serum measurements and the goal is to determine which of these are relevant for forecasting diabetes progression.
We fit a neural network with one hidden layer $L=1$ and $D_{H}=20$ and perform the node selection algorithm with the FDR bounded by $\alpha=0.01$. Figure 8 illustrates the shrinkage and node selection and compares the coefficients of the Lasso linear model with cross-validation (LassoCV) to the original and the sparsified weights of our model. Predictors with considerable effect obtained by both models coincide, whilst some of the variables the Lasso model excludes (e.g. age) are still present in the VBNN’s estimates. Compared with Lasso, VBNN has the advtange of learning potential nonlinear relationships between disease progression and the predictors, which is explored in Figure 9, illustrating the predictive means and uncertainty of the observations of VBNN for four of the predictors (with all other predictors are fixed to their mean) compared to the predictions of Lasso. While the uncertainty is wide, the results suggest potential nonlinear relationships, e.g. with lamotrigine and age, the latter of which is not selected in Lasso.
Moreover, Figure 9 highlights how predictions obtained from the sparse version of the variational predictive distribution almost overlap, thus providing a reasonable, cheaper approximation. However, we note that predictive performance is only slightly improved with VBNN (see Table 2 and supplementary Figure 10 in Section C.3).
Table 2: RMSE, NLL and coverage for diabetes dataset.
@@
LassoCV $53.7\pm.5$ $5.4\pm.13$$96\pm.03$
VBNN $52.9\pm 4$ $5.4\pm.09$$96\pm.03$
SVI $55.\pm 5$ $6.5\pm.5$$63\pm.07$
BBB $54.2\pm.1$ $5.4\pm.08$$94\pm.02$
Figure 8: Coefficients of LassoCV regression (on the left), predictive means of the weights of the neural network (in the middle) and predictive means of the sparse weights obtained for $\alpha=0.01$ (on the right).
Figure 9: Slices of the predictive mean and uncertainty for observations for four predictors obtained by VBNN with and without node selection and by Lasso with cross-validation.

### v3

@@ 3.4 Inferring the Network Structure
The choice of the network architecture has significant practical implications on the generalization of the model, and so sparsity-promoting priors for the network weights have emerged as a promising approach for increased robustness to the choice of architecture. In this article, we focus on a class of continuous shrinkage priors, which results in more tractable computations, yet also implies non-zero posterior means and does not lead to automatic network architecture selection. To address this, we consider a post-processing node selection algorithm to obtain a sparse variational posterior, which both aids interpretability and improves scalability of predictive inference (discussed in Section 3.5)
In Bayesian linear regression with shrinkage priors, several post-processing methods have been proposed to yield a sparse solution (see e.g.(Piironen et al., 2020; Li and Pati, 2017; Griffin, 2024)). The method known as decoupling shrinkage and selection (DSS) (Hahn and Carvalho, 2015) obtains sparse estimates of the weights by minimizing the sum of the predictive loss function with a parsimony-inducing penalty. An alternative approach is the the penalized credible regions (PenCR) method (Zhang et al., 2021), which identifies the "sparsest" solution in posterior credible regions corresponding to different levels; it is shown to perform well in the case of global-local shrinkage priors and under certain assumptions, PenCR produces the same results as DSS. Similarly, we propose to make use of credible intervals to select nodes. Following (Li and Lin, 2010), we implement an automatic credible interval criterion which selects a node as long as its credible interval doesn’t cover zero. Specifically, recall that the variational posterior of the weights is $W_{l,d,d^{\prime}}\sim\text{N}(m^{W}_{l,d,d^{\prime}},B^{W}_{l,d,d^{\prime}})$ for $l=1,\ldots,L,d=1,\dots,D_{l},d^{\prime}=1,\ldots,D_{l-1}$, where $\mathbf{m}^{W}_{l,d}$ and $\mathbf{B}^{W}_{l,d}$ denote the subsets of the mean $\mathbf{m}_{l,d}$ and covariance matrix $\mathbf{B}_{l,d}$ corresponding to the weights. Then, we obtain sparse weights $\widehat{W}_{l,d,d^{\prime}}$ with sparse variational distribution $\widehat{q}(b_{l,d},\widehat{\mathbf{W}}_{l,d})$ for some $l\in\mathcal{L}\subseteq\{1,\ldots,L+1\},\;d\in\mathcal{D}_{l}\subseteq\{1,\ldots,D_{l}\},d^{\prime}\in\mathcal{D}_{l-1}\subseteq\{1,\ldots,D_{l-1}\}$, defined by setting:
@@
where $Q(W_{l,d,d^{\prime}}<0)=1-Q(W_{l,d,d^{\prime}}>0)=\Phi(-m^{W}_{l,d,d^{\prime}}/\sqrt{(B^{W}_{l,d})_{d^{\prime}d^{\prime}}})$. The threshold $\kappa$ is chosen to control the Bayesian false discovery rate, which is calculated as
@@
Once we sweep through $\mathcal{Q}_{l,d.d^{\prime}}$, we do a backwards pass to remove the nodes with no connections. If the node has no outgoing connections, then all the incoming connections need to be removed, and conversely, if the node has no incoming connections, then all the outgoing connections can be removed. An example of the network resulting after applying the Algorithm 2 is illustrated by the Figure 7.
@@ 3.5 Predictions
Thus, before computing predictions, we first iterate to update the variational predictive distributions of $\mathbf{a}_{*}$, $\bm{\gamma}_{*}$, and $\bm{\omega}_{*}$. The corresponding ELBO (derived in the Section B.2) is monitored for convergence. While a closed-form expression for the integral in Equation 22 is unavailable, generating samples from the variational predictive is straightforward;
@@
We can also compute the expectation and variance of $\mathbf{y}_{*}$ in closed form. Specifically, the expectation of $\mathbf{y}_{*}$ under the variational predictive distribution is:
@@
Sparse prediction
Observe that the variational algorithm used for prediction scales linearly with the number of hidden layers and the number of samples but cubically with the number of hidden units, this motivates the node-selection method proposed in Section 3.4. To save on both computation and storage, the variational predictive distribution can be computed based on the sparse variational posterior (Section 3.4). For a new data point $\mathbf{x}_{*}$, we obtain expectation and variance of $\mathbf{y}_{*}$ by first computing the sparse versions of variational predictive distributions $\widehat{q}(\mathbf{a}_{*}),\widehat{q}(\bm{\gamma}_{*})$ and $\widehat{q}(\bm{\omega}_{*})$ as in Equations 24, 19 and 23 by plugging $\widehat{q}(b_{l,d},\widehat{\mathbf{W}}_{l,d})$ instead of the $q(b_{l,d},\mathbf{W}_{l,d})$, which only requires updates for the subset of nodes with nonzero weights.
@@ 3.6 Ensembles of Variational Approximations
While the variational algorithm described in Section 3.1 increases the ELBO at each epoch, the ELBO is a non-convex function of the variational parameters and only convergence to a local optimum is guaranteed. Due to identifiability issues, the posterior distribution of a Bayesian neural network is highly multimodal, and exploring this posterior is notoriously challenging (Papamarkou et al., 2022). A single variational approximation tends to concentrate around one mode and can understate posterior uncertainty. Several approaches have been proposed to overcome such issues. Recently, (Ohn and Lin, 2024) introduced adaptive variational inference, which achieves optimal posterior contraction rate and model selection consistency by considering several variational approximations obtained in different models. Similarly, (Yao et al., 2022) introduced an approach which uses parallel runs of inference algorithms to cover as many modes of the posterior distribution as possible and then combines these using Bayesian stacking. While the idea of combining the outputs of several neural networks is not novel (Hansen and Salamon, 1990; Levin et al., 1990), we note that deep non-Bayesian ensembles recently received quite a lot of attention (Lakshminarayanan et al., 2017) and several alternatives and modifications enabling uncertainty quantification were proposed (Wilson and Izmailov, 2020; D' Angelo and Fortuin, 2021; Pearce et al., 2020; Wenzel et al., 2020b). Observing non-optimality of conventional deep ensembles combining point estimates (Wu and Williamson, 2024), in our approach, we adopt the ideas of (Ohn and Lin, 2024; Yao et al., 2022) and consider ensembles of posterior approximations in a similar but simpler fashion. Specifically, we consider an ensemble of variational approximations, obtained by running in parallel the variational algorithm multiple times with different random starting points and combining those with respect to the optimization objective. In this case, letting $k=1,\ldots,K$ index the different variational approximations, we compute the weight $w_{k}$ associated with each approximation in accordance with the tempered ELBO:
$w_{k}\propto\exp\left(\zeta\text{ELBO}_{k}\right),$
where $\zeta$ is a tempering parameter, setting which to be in the interval $(0,0.1]$ allows for avoiding a strong dominance of a single particular model. We can interpret this as a Bayesian model averaging (BMA) across the $K$ different models/approximations. While in a classical BMA setting, the weights would be proportional to the marginal likelihood for each model, the use of the ELBO is motivated as it provides a lower bound to the marginal likelihood and can be computed in closed form. Next, we compute predictions by taking a weighted average of the predictive distributions of each model (given in Equation 26), that is
@@ 4 Experiments
We evaluate the variational bow tie neural network (VBNN) on several datasets. First, we consider a simple nonlinear synthetic example to compare with a ground truth. We then validate VBNN on the diabetes dataset, first considered in (Efron et al., 2004) to demonstrate the least angle regression (LARS) algorithm for variable selection, and subsequently, used in different proposals for sparsity-promoting priors algorithms (e.g. (Park and Casella, 2008; Li and Lin, 2010)). Lastly, we consider a range of popular regression datasets from the UCI Machine Learning Repository (M. et al., 2007).
The importance of suitable initialization choice in NNs is well known (Wenzel et al., 2020a; Daniely et al., 2016; He et al., 2015), and we design two possible random initialization schemes of the VBNN, which are described in Section D.1 and used in all experiments. Convergence of the ELBO is monitored during the training and prediction stages, where if three consecutive measurements of ELBO for training differ by less than the specified threshold, the phase is stopped and the model moves to the prediction stage, where we proceed similarly. In most experiments, the thresholds during the training and prediction stages are set, respectively, to $51e-$ 5 and $41e-$ 4.
We compare the performance of VBNN to various frameworks (summarized in Table 2), namely, to BNNs inferred with automatic differentiation VI with the mean-field variational family (mfVI) (Kucukelbir et al., 2017) and Hamiltonian Monte Carlo with the No-U-Turn sampler (HMC) (Hoffman et al., 2014) implemented in Numpyro (Phan et al., 2019), Bayes by Backprop (BBB) (Blundell et al., 2015) implemented with Pytorch, and BNNs with Horseshoe priors (HSBNN) of (Ghosh et al., 2019) which considers a structured variational family and learns the variational parameters by obtaining gradient estimators. We also consider the variational bow tie neural network with Gaussian priors (GVBNN), in contrast to shrinkage priors. As CAVI may be expensive for large data, in Figure 10 we additionally consider VBNN inferred with SVI (SVBNN). Further comparisons between the performance of VBNN and SVBNN are provided in Section D.6.
Table 2: List of the models considered to evaluate the performance of our method.
Model Description
mfVI BNN with Automatic Differentiation VI with mean-field family
HMC BNN with Hamiltonian Monte Carlo with No-U-Turn sampler
BBB Bayes by Backprob
HSBNN BNN with Horseshoe priors inferred with Black Box VI
VBNN Our model inferred with CAVI
GVBNN VBNN with Gaussian priors inferred with CAVI
SVBNN Our model inferred with SVI
For all the datasets, we evaluate the performance over 10 random splits, where we use $90\%$ of the data for the training and $10\%$ for testing the model. We record the root mean squared error (RMSE), the predictive negative log-likelihood of the test data (NLL) and the empirical coverage (EC) (see Appendix D for additional implementation details). The empirical coverage reports the fraction of observations contained within the $(1-\alpha)*100\%$ credible intervals (CIs), which are computed based on the Gaussian approximation. In the ideal setting, the computed EC should equal the CI level. More specifically, let $\mathbf{y}^{*}_{i}$ be the true target for test points $i=1,\ldots,N^{*}$ and $q_{\alpha/2}$ and $q_{1-\alpha/2}$ be the quantiles based on the model’s Gaussian approximation for some $\alpha\in[0.5,1]$, then
$\text{EC}(\alpha)$ $=\frac{1}{N^{*}}\sum_{i=1}^{N^{*}}\textbf{1}(\mathbf{y}^{*}_{i}\in[q_{\alpha/2},q_{1-\alpha/2}]).$
If $\text{EC}(\alpha)>1-\alpha$ then the CIs are too wide; a worse scenario occurs when $\text{EC}(\alpha)<1-\alpha$ as it means that the CIs are too narrow and the model is overconfident in its predictions.
@@ 4.1 Simulated Example
We construct a synthetic dataset generated by first uniformly sampling a two-dimensional input vector $\mathbf{x}_{n}=(x_{n,1},x_{n,2})$, with $x_{n,d}\sim\text{Unif}([-2,2])$, and assume only the first feature influences the output: $y_{n}=f(x_{n,1})+\epsilon_{n}=0.1x_{n,1}^{2}+10\sin(x_{n,1})+\epsilon_{n}$, where $\epsilon_{n}\sim\text{N}(0,0.5)$. Then, the dataset consisting of $N=300$ observations is used to investigate the performance of VBNN compared to the GVBNN, mfVI, HMC, BBB and HSBNN baselines as we increase the number of hidden layers, setting $L=1,2$ or $4$, whilst keeping the number of hidden units in each layer fixed to $D_{H}=20$. In general, for this simple non-linear example, the performance tends to deteriorate with increasing architecture complexity (larger depth). While HMC performs consistently well across all depths, the cost associated with the sampling approach is high. VBNN is competitive to HSBNN and outperforms mfVI, GVBNN and BBB in terms of accuracy (see Figure 5). Further, except for HMC, the empirical coverage of VBNN is the most robust to the choice of depth; for the largest choice of $L=4$, mfVI, BBB, GVBNN and even HSBNN provide overly wide CIs while VBNN more closely reaches the desired coverage (see Figure 6).
Figure 5: Simulated example. Performance in terms of the RMSE and NLL as the depth increases for different models and algorithms. HMC can be seen as a gold standard. VBNN is competitive with HSBNN and is more robust to the choice of depth and overparameterization than GVBNN, mfVI, BBB.
Figure 6: Simulated example. Empirical coverage (which is the fraction of observations contained within the CIs of level $1-\alpha$) as a function of CI level for the simulated dataset for three different settings of the network’s depth. The dashed gray line depicts the ideal scenario with empirical coverage equal to CI level, while above and below the gray line indicate coverage greater or less than CI level, i.e. CIs are too wide or too small, respectively.
@@
For each depth, Figure 7 illustrates the predictive means and uncertainties computed for the observations as well as DAGs of networks’ structures obtained after the post-process node selection algorithm described in Section 3.4, where the Bayesian false discovery rate is constrained by setting the error rate to $\alpha=0.01$. The sparsity-promoting prior combined with the node selection algorithm can effectively prune the over-parametrized neural networks; for example, the sparse one-layer neural network contains only $11$ hidden nodes with $33$ total edges/weights from the initial $D_{H}=20$ with $60$ total edges/weights. Moreover, the estimated regression function and credible intervals both from the variational predictive and the sparse variation predictive distribution, recover the true function well. In this way, VBNN provides an effective tool to reduce predictive computational complexity and storage as well as ease interpretation. Note that the predictions show no relation with the coordinate $x_{2}$ (Figure 7(a)), recovering the true function, but some of the connections from $x_{2}$ are still present in the sparse network (Figure 7(c)), due to identifiability issues, although with overall low weight.
(a) The predictive mean as a function of the second coordinate for $L=1$.
(b) The predictive mean and uncertainty quantification for the observations for different depth
(c) The architecture of the network for the bound on the FDR $\alpha=0.01$ for depths $L=1,2,4$ (left to right).
@@
Figure 7: Simulated example. Predictive means and pointwise CIs computed for the observations as a function of the second coordinate (a) and first coordinate for different depths (b). The architecture of the network is visualized in (c) for the bound on the FDR $\alpha=0.01$ for different settings of the network’s depth of $L=1,2,4$ (left to right).
@@ 4.2 Diabetes Example
The diabetes data consists of $n=442$ entries obtained for $p=10$ input variables and a quantitative response measuring disease progression. The predictors are age, sex, body mass index, average blood pressure and six blood serum measurements and the goal is to determine which of these are relevant for forecasting diabetes progression. We fit a neural network with one hidden layer $L=1$ and $D_{H}=20$ and perform the node selection algorithm with the FDR bounded by $\alpha=0.01$. Figure 8 illustrates the shrinkage and node selection algorithm and compares the coefficients of the Lasso linear model (Tibshirani, 1996) to the original and the sparsified weights of our model. Lasso regression produces sparse coefficients by minimising the residual sum of squares with an added penalty term; the penalty parameter crucially determines the level of sparsity and is tuned with cross-validation (LassoCV). Predictors with considerable effect obtained by both models coincide, whilst some of the variables the Lasso model excludes (e.g. age) are still present in the VBNN’s estimates. Compared with Lasso, VBNN has the advantage of learning potential nonlinear relationships between disease progression and the predictors, which is explored in Figure 9, illustrating the predictive means and uncertainty of the observations of VBNN for four of the predictors (with all other predictors are fixed to their mean). While the uncertainty is wide, the results suggest potential nonlinear relationships, e.g. with lamotrigine and age, the latter of which is not selected in Lasso. Moreover, Figure 9 highlights how predictions obtained from the sparse version of the variational predictive distribution almost overlap, thus providing a reasonable, cheaper approximation. However, we note that the predictive performance is similar to LassoCV, with the most competitive methods being VBNN, HMC and BBB (see Table 3 and supplementary Figure 12 in Section D.3).
Table 3: RMSE, NLL and empirical coverage for diabetes dataset.
@@
LassoCV $54.2\pm 6.5$ $5.4\pm.13$$96\pm.03$
mfVI $57.2\pm 7.4$ $9.3\pm 1.7$$47\pm.1$
HMC $54.5\pm 7.8$ $5.4\pm.16$$96\pm.04$
BBB $54.9\pm 7.3$ $5.49\pm.18$$94\pm.04$
HSBNN $56.8\pm 7.4$ $6.8\pm 1.6$$67\pm.15$
GVBNN $55.65\pm 7.8$ $5$14 .$96\pm.04$
VBNN $54.5\pm 7.2$ $5.4\pm.15$$96\pm.04$
Figure 8: Diabetes example. Coefficients of LassoCV regression (on the left), posterior means of the weights of the neural network (in the middle) and posterior means of the sparse weights obtained for $\alpha=0.01$ (on the right). For illustrative purposes, absolute values of the coefficients and weights are shown with max-min scaling.
Figure 9: Diabetes example. Slices of the predictive mean and pointwise credible intervals for observations as a function of four predictors obtained by VBNN with and without node selection and by Lasso with cross-validation.